# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ 5.4 Comparison to the ODIN Survey
LATIS1-D2-2 | 149.6910 | 2.1750 | 2.679 | 149.6475 | 2.1254 | 2.711 | 179 | 1.94 | 2997, 3262 | 15 |
LATIS1-D2-3 | 150.3470 | 2.2740 | 2.457 | 150.3403 | 2.2236 | 2.496 | 370 | 1.53+ | 1982 | 15 |
LATIS1-D2-4 | 150.0330 | 2.1750 | 2.685 | 150.0188 | 2.2508 | 2.658 | 105 | 2.27+ | 2801 | 15 |
@@
C.PS14 | 150.1900 | 2.2900 | 2.816 | 150.1916 | 2.2928 | 2.805 | 487 | 0.09 | 3619 | 1,2 |
C.PS15 | 149.9480 | 2.3070 | 2.816 | 149.9523 | 2.3208 | 2.830 | 361 | 0.42 | 3780 | 1,2 |
PClJ1000+0200 | 150.0910 | 1.9920 | 2.911 | 150.0675 | 1.9907 | 2.895 | 1056 | 0.68+ | 4228 | 17 |
C.PS17 | 149.9110 | 2.2680 | 2.921 | 149.9077 | 2.1100 | 2.924 | 507 | 4.56∗ | 4401 | 1,2 |
C.PS18 | 150.2500 | 2.2790 | 2.953 | 150.2746 | 2.2719 | 2.951 | 1418 | 0.74 | 4375, 4596‡ | 1,2 |
C.PS19 | 149.7000 | 1.9790 | 3.002 | 149.9725 | 2.3217 | 3.004 | 181 | 0.59 | 4681 | 1,2 |
C.PS20 | 149.9980 | 2.3000 | 3.035 | 149.9725 | 2.3217 | 3.035 | 235 | 0.95 | 4795 | 1,2 |
MAGAZ3NEJ1001 | 150.4417 | 2.5192 | 3.123 | 150.4869 | 2.5157 | 3.141 | 118 | 1.28 | 4479‡ | 18 |
C.PS21 | 149.6900 | 1.8090 | 3.232 | 149.6874 | 1.8094 | 3.232 | 63 | 0.08 | 5341 | 1,2 |
C.PS22 | 149.8810 | 1.8500 | 3.232 | 149.8937 | 1.8543 | 3.238 | 246 | 0.38 | 5378 | 1,2 |
Elentári1 | 149.8060 | 2.4510 | 3.366 | 149.9001 | 2.3145 | 3.353 | 931 | 4.58∗ | 5400‡, 5713 | 19, 20 |
Elentári2 | 149.8774 | 2.2850 | 3.341 | 149.9001 | 2.3145 | 3.353 | 931 | 1.03 | 5400‡, 5713 | 19 |
Elentári3 | 149.9369 | 2.2749 | 3.269 | 149.8863 | 2.3472 | 3.282 | 159 | 2.46+ | 5400‡ | 19 |
Elentári4 | 150.4021 | 2.3356 | 3.255 | 150.4004 | 2.3382 | 3.258 | 116 | 0.09 | 5401 | 19 |
Elentári5 | 150.2153 | 2.4941 | 3.343 | 150.2141 | 2.4612 | 3.329 | 231 | 0.91 | 5598 | 19 |
Elentári6 | 149.8287 | 2.4306 | 3.315 | 149.9001 | 2.3145 | 3.353 | 931 | 3.78∗,+ | 5400‡, 5713 | 19 |
Elentári7 | 150.1099 | 2.5634 | 3.354 | 150.2141 | 2.4612 | 3.329 | 231 | 4.03∗ | 5598 | 19, 20 |
Elentári8 | 149.8869 | 2.3242 | 3.248 | 149.8863 | 2.3472 | 3.282 | 159 | 0.64 | 5400‡ | 19 |
@@
Taralay1 | 150.3522 | 2.3540 | 4.567 | 150.3551 | 2.3534 | 4.573 | 1179 | 0.07 | 7353 | 2,21 |
Taralay2 | 150.1771 | 2.3013 | 4.592 | 150.1372 | 2.3240 | 4.609 | 450 | 1.12 | 7385 | 2,21 |
References. — [1] Lemaux et al. (2014a); [2] Lemaux et al. (2018); [3] Ata et al. (2022); [4] Yuan et al. (2014); [5] Koyama et al. (2021); [6] Polletta et al. (2021); [7] Darvish et al. (2020); [8] Franck & McGaugh (2016); [9] Cucciati et al. (2018); [10] Casey et al. (2015); [11] Diener et al. (2013); [12] Lee et al. (2016); [13] Chiang et al. (2015); [14] Wang et al. (2016); [ 15] Newman et al. (2022); [16] Ito et al. (2023); [17] Cucciati et al. (2014); [18] McConachie et al. (2024); [19] Forrest et al. (2023); [20] McConachie et al. (2022); [21] Staab et al. (2024); [22] Kakimoto et al. (2023)
@@
Structure | RAlit | Declit | RAC3VO | DecC3VO | Amp. | dtrans | IDC3VO | ref. | ||
@@
VANDELS1 | 53.0629 | -27.7237 | 2.290 | 53.0742 | -27.7048 | 2.304 | 1131 | 0.65 | 278 | 1 |
VANDELS2 | 53.1412 | -27.6871 | 2.300 | 53.0742 | -27.7048 | 2.304 | 1131 | 1.88 | 278 | 1 |
VANDELS3 | 53.1421 | -27.8221 | 2.340 | 53.0742 | -27.7048 | 2.304 | 1131 | 4.00+ | 278 | 1 |
E.PS1 | 53.0925 | -27.7158 | 2.451 | 53.0570 | -27.6994 | 2.442 | 606 | 1.07 | 416, 471 | 2,3 |
E.PS2 | 53.1348 | -27.9125 | 2.559 | 53.0828 | 27.9125 | 2.569 | 490 | 1.56+ | 598 | 2,3 |
E.PS3 | 53.1504 | -27.8098 | 2.622 | 53.0631 | -27.7850 | 2.613 | 265 | 2.40+ | 531 | 2,3 |
Drishti1 | 53.0731 | -27.9323 | 2.674 | 53.0690 | -27.8883 | 2.692 | 715 | 1.30 | 762 | 4 |
Drishti2 | 53.1876 | -27.7943 | 2.694 | 53.2178 | -27.7836 | 2.699 | 431 | 0.85 | 768 | 4 |
Drishti3 | 53.1133 | -27.8984 | 2.697 | 53.0690 | -27.8883 | 2.692 | 715 | 1.19 | 762 | 4 |
Surabhi | 52.9988 | -27.8063 | 2.795 | 52.9321 | -27.8681 | 2.774 | 186 | 2.49 | 848 | 4 |
VANDELS6 | 53.2037 | -27.7746 | 2.800 | 53.0948 | -27.8190 | 2.813 | 1279 | 3.09∗ | 880 | 1 |
VANDELS7 | 53.1412 | -27.8612 | 3.170 | 53.1955 | -27.7507 | 3.189 | 750 | 3.39∗,+ | 1102 | 1 |
VANDELS8 | 53.1346 | -27.6954 | 3.230 | 53.1955 | -27.7507 | 3.189 | 750 | 2.16 | 1102 | 1 |
@@
Shrawan1 | 53.2727 | -27.7936 | 3.343 | 53.2721 | -27.7902 | 3.344 | 288 | 0.09 | 1224 | 4 |
Shrawan2 | 53.0714 | -27.9353 | 3.355 | 53.0346 | -27.9438 | 3.363 | 128 | 0.93 | 1246 | 4 |
Shrawan3 | 53.1552 | -27.8959 | 3.242 | 53.1510 | -27.8919 | 3.237 | 342 | 0.15 | 1129 | 4 |
Shrawan4 | 53.2022 | -27.9406 | 3.335 | 53.1984 | -27.9349 | 3.337 | 186 | 0.18 | 1222 | 4 |
Smruti1 | 53.0076 | -27.7463 | 3.410 | 53.0193 | -27.7403 | 3.417 | 286 | 0.33 | 1270 | 4,5,6,7 |
Smruti2 | 53.0042 | -27.7411 | 3.479 | 53.0159 | -27.7498 | 3.490 | 234 | 0.37 | 1289 | 4,5,6,7 |
Smruti3 | 53.0613 | -27.8723 | 3.471 | 53.8723 | -27.8795 | 3.469 | 1731 | 1.96+ | 1308 | 4,5,6,7 |
Smruti4 | 53.2290 | -27.8828 | 3.462 | 53.8723 | -27.8795 | 3.469 | 1731 | 2.11+ | 1308 | 4,5,6,7 |
Smruti5 | 53.0412 | -27.7804 | 3.530 | 53.0159 | -27.7498 | 3.490 | 234 | 1.03 | 1289 | 4,5,6,7 |
Smruti6 | 53.1586 | -27.6964 | 3.418 | 53.1521 | -27.7078 | 3.418 | 121 | 0.25+ | 1266 | 4,5,6,7 |
Sparsh | 53.0579 | -27.8670 | 3.696 | 53.0478 | -27.8730 | 3.701 | 1173 | 0.29 | 1378 | 4,8 |
Ruchi1 | 53.2124 | -27.8306 | 4.150 | 53.2185 | -27.8273 | 4.155 | 610 | 0.16 | 1422 | 4 |
@@
Structure | RAlit | Declit | RAC3VO | DecC3VO | Amp. | dtrans | IDC3VO | ref. | ||
@@
PClJ0227-0421.1 | 36.6375 | -4.2381 | 3.269 | 36.6969 | -4.3563 | 3.292 | 1574 | 3.69∗ | 2820 | 1,5 |
PClJ0227-0421.2 | 36.7438 | -4.2949 | 3.325 | 36.6969 | -4.3563 | 3.292 | 1574 | 2.14+ | 2820 | 1,5 |
PClJ0227-0421.3 | 36.7775 | -4.3457 | 3.299 | 36.6969 | -4.3563 | 3.292 | 1574 | 2.25+ | 2820 | 1,5 |
PClJ0227-0421.4 | 36.6467 | -4.3703 | 3.300 | 36.6969 | -4.3563 | 3.292 | 1574 | 1.44 | 2820 | 1,5 |
PClJ0227-0421.5 | 36.5817 | -4.3877 | 3.298 | 36.6969 | -4.3563 | 3.292 | 1574 | 3.31∗,+ | 2820 | 1,5 |
PClJ0227-0421.6 | 36.6417 | -4.4472 | 3.302 | 36.6434 | -4.4638 | 3.258 | 75 | 0.47 | 2773 | 1,5 |
@@ 6 Conclusions
By applying our search algorithm to our plethora of photometry and spectroscopy in the C3VO fields, we were able to find detections of 565 protostructure candidates, many of which have not previously been reported in the literature. The main takeaways from our findings are:
@@
Over 400 potentially new protostructures were found in the COSMOS, ECDFS, and CFHTLS-D1 fields, all of which have some spectroscopically-confirmed members.
With the aid of mock observations of simulated structures, we have found completeness and purity as high as 80% and 70%, respectively, in the best-case scenario. These numbers correspond to and structures with for completeness, and a spectral redshift fraction of % for purity (i.e., SF1.5 or better). Typical completeness and purity are around 20-40% for , though the precise value sensitively depends on the mass and redshift, with a precipitous falloff to . Purity is typically found to be 40-60%.
More than 90% of the 100 protoclusters previously reported in the literature (which were originally selected using a variety of different methods) in these three fields were recovered with our method. Of the few protoclusters that were not detected using our methodology, the vast majority were either situated near the edge of our maps or in multi-component systems where our deblending was not aggressive enough to discriminate different sub-components.
@@
The RA/Dec/z coordinates of all protostructure candidates are made publicly available as an electronic table on CDS. The Gaussian amplitude () of each protostructure candidate, which represents a rough mass proxy, as well as its associated error, is also provided.
Fits are provided between our mass proxy () and the mass of simulated structures for differing levels of spectroscopic completeness. These fits may be applied to the Gaussian amplitude of our candidate protostructure detections to derive crude masses for the observed candidates, though with considerable scatter and some bias. We caution that masses of our candidates can be estimated more accurately and precisely, but it must be on a case-by-case basis considering individual parameters such as the spectral sampling, tracer population, and galaxy bias for the tracers as has been done with other C3VO studies (e.g., Cucciati et al., 2018; Shen et al., 2021; Forrest et al., 2023).
With our results, we have once again demonstrated the efficacy of the VMC mapping method and its utility when combined with any large and complete spectroscopic and photometric data set. While our work was on a fairly small region of the sky, the methodology is adaptable to any spectroscopic and photometric data set of sufficient size. This is especially important to consider in the context of active and upcoming large-scale surveys, such as those conducted by Euclid121212https://www.euclid-ec.org/, the Vera C. Rubin Observatory131313https://rubinobservatory.org/, Subaru’s Prime Focus Spectrograph (PFS)141414https://pfs.ipmu.jp, the Nancy Grace Roman Space Telescope151515https://roman.gsfc.nasa.gov/, and the Maunakea Spectroscopic Explorer (MSE)161616https://mse.cfht.hawaii.edu/, which are able to reach deeper magnitudes and offer larger data sets at even higher redshifts than ever before. Of these, Euclid is the only telescope currently in operation, and the Euclid Wide Survey in particular has the potential to find tens of thousands of protoclusters with masses over in its 14,500 deg2 area of coverage (Euclid Collaboration et al., 2024). Methods based largely on the VMC technique are already proving to be effective at uncovering a large number of protocluster candidates in EUCLID mock data and will be used for searches in real data (Ramos-Chernenko in prep). Thanks to the deluge of data on the way focused on the redshift range that marks the transition between clusters and protoclusters, the next decade promises to be a boon for protocluster science.

### v2

@@ 5.4 Comparison to the ODIN Survey
LATIS1-D2-2 | 149.6910 | 2.1750 | 2.679 | 149.6475 | 2.1254 | 2.711 | 179 | 1.94 | C_SP137 | 15 |
LATIS1-D2-3 | 150.3470 | 2.2740 | 2.457 | 150.3403 | 2.2236 | 2.496 | 370 | 1.53+ | C_SP22 | 15 |
LATIS1-D2-4 | 150.0330 | 2.1750 | 2.685 | 150.0188 | 2.2508 | 2.658 | 105 | 2.27+ | C_SP165 | 15 |
@@
C.PS14 | 150.1900 | 2.2900 | 2.816 | 150.1916 | 2.2928 | 2.805 | 487 | 0.09 | C_SP12 | 1,2 |
C.PS15 | 149.9480 | 2.3070 | 2.816 | 149.9523 | 2.3208 | 2.830 | 361 | 0.42 | C_SP23 | 1,2 |
PClJ1000+0200 | 150.0910 | 1.9920 | 2.911 | 150.0675 | 1.9907 | 2.895 | 1056 | 0.68+ | C_SP4 | 17 |
C.PS17 | 149.9110 | 2.2680 | 2.921 | 149.9077 | 2.1100 | 2.924 | 507 | 4.56∗ | C_SP11 | 1,2 |
C.PS18 | 150.2500 | 2.2790 | 2.953 | 150.2746 | 2.2719 | 2.951 | 1418 | 0.74 | C_MP2‡ | 1,2 |
C.PS19 | 149.7000 | 1.9790 | 3.002 | 149.9725 | 2.3217 | 3.004 | 181 | 0.59 | C_SP80 | 1,2 |
C.PS20 | 149.9980 | 2.3000 | 3.035 | 149.9725 | 2.3217 | 3.035 | 235 | 0.95 | C_SP55 | 1,2 |
MAGAZ3NEJ1001 | 150.4417 | 2.5192 | 3.123 | 150.4869 | 2.5157 | 3.141 | 118 | 1.28 | C_MP52‡ | 18 |
C.PS21 | 149.6900 | 1.8090 | 3.232 | 149.6874 | 1.8094 | 3.232 | 63 | 0.08 | C_SP223 | 1,2 |
C.PS22 | 149.8810 | 1.8500 | 3.232 | 149.8937 | 1.8543 | 3.238 | 246 | 0.38 | C_SP47 | 1,2 |
Elentári1 | 149.8060 | 2.4510 | 3.366 | 149.9001 | 2.3145 | 3.353 | 931 | 4.58∗ | C_MP4‡ | 19, 20 |
Elentári2 | 149.8774 | 2.2850 | 3.341 | 149.9001 | 2.3145 | 3.353 | 931 | 1.03 | C_MP4‡ | 19 |
Elentári3 | 149.9369 | 2.2749 | 3.269 | 149.8863 | 2.3472 | 3.282 | 159 | 2.46+ | C_MP46‡ | 19 |
Elentári4 | 150.4021 | 2.3356 | 3.255 | 150.4004 | 2.3382 | 3.258 | 116 | 0.09 | C_SP148 | 19 |
Elentári5 | 150.2153 | 2.4941 | 3.343 | 150.2141 | 2.4612 | 3.329 | 231 | 0.91 | C_SP57 | 19 |
Elentári6 | 149.8287 | 2.4306 | 3.315 | 149.9001 | 2.3145 | 3.353 | 931 | 3.78∗,+ | C_MP4‡ | 19 |
Elentári7 | 150.1099 | 2.5634 | 3.354 | 150.2141 | 2.4612 | 3.329 | 231 | 4.03∗ | C_SP57 | 19, 20 |
Elentári8 | 149.8869 | 2.3242 | 3.248 | 149.8863 | 2.3472 | 3.282 | 159 | 0.64 | C_MP46‡ | 19 |
@@
Taralay1 | 150.3522 | 2.3540 | 4.567 | 150.3551 | 2.3534 | 4.573 | 1179 | 0.07 | C_SP3 | 2,21 |
Taralay2 | 150.1771 | 2.3013 | 4.592 | 150.1372 | 2.3240 | 4.609 | 450 | 1.12 | C_SP15 | 2,21 |
References. — [1] Lemaux et al. (2014a); [2] Lemaux et al. (2018); [3] Ata et al. (2022); [4] Yuan et al. (2014); [5] Koyama et al. (2021); [6] Polletta et al. (2021); [7] Darvish et al. (2020); [8] Franck & McGaugh (2016); [9] Cucciati et al. (2018); [10] Casey et al. (2015); [11] Diener et al. (2013); [12] Lee et al. (2016); [13] Chiang et al. (2015); [14] Wang et al. (2016); [ 15] Newman et al. (2022); [16] Ito et al. (2023); [17] Cucciati et al. (2014); [18] McConachie et al. (2025); [19] Forrest et al. (2023); [20] McConachie et al. (2022); [21] Staab et al. (2024); [22] Kakimoto et al. (2024).
@@
Structure | R.A.lit | Decl.lit | R.A.C3VO | Decl.C3VO | Amp.a | dtrans | IDC3VOb | ref. | ||
@@
VANDELS1 | 53.0629 | -27.7237 | 2.290 | 53.0742 | -27.7048 | 2.304 | 1131 | 0.65 | E_SP5 | 1 |
VANDELS2 | 53.1412 | -27.6871 | 2.300 | 53.0742 | -27.7048 | 2.304 | 1131 | 1.88 | E_SP5 | 1 |
VANDELS3 | 53.1421 | -27.8221 | 2.340 | 53.0742 | -27.7048 | 2.304 | 1131 | 4.00+ | E_SP5 | 1 |
E.PS1 | 53.0925 | -27.7158 | 2.451 | 53.0570 | -27.6994 | 2.442 | 606 | 1.07 | E_SP20 | 2,3 |
E.PS2 | 53.1348 | -27.9125 | 2.559 | 53.0828 | 27.9125 | 2.569 | 490 | 1.56+ | E_SP13 | 2,3 |
E.PS3 | 53.1504 | -27.8098 | 2.622 | 53.0631 | -27.7850 | 2.613 | 265 | 2.40+ | E_SP39 | 2,3 |
Drishti1 | 53.0731 | -27.9323 | 2.674 | 53.0690 | -27.8883 | 2.692 | 715 | 1.30 | E_SP8 | 4 |
Drishti2 | 53.1876 | -27.7943 | 2.694 | 53.2178 | -27.7836 | 2.699 | 431 | 0.85 | E_SP18 | 4 |
Drishti3 | 53.1133 | -27.8984 | 2.697 | 53.0690 | -27.8883 | 2.692 | 715 | 1.19 | E_SP8 | 4 |
Surabhi | 52.9988 | -27.8063 | 2.795 | 52.9321 | -27.8681 | 2.774 | 186 | 2.49 | E_SP45 | 4 |
VANDELS6 | 53.2037 | -27.7746 | 2.800 | 53.0948 | -27.8190 | 2.813 | 1279 | 3.09∗ | E_SP3 | 1 |
VANDELS7 | 53.1412 | -27.8612 | 3.170 | 53.1955 | -27.7507 | 3.189 | 750 | 3.39∗,+ | E_SP7 | 1 |
VANDELS8 | 53.1346 | -27.6954 | 3.230 | 53.1955 | -27.7507 | 3.189 | 750 | 2.16 | E_SP7 | 1 |
@@
Shrawan1 | 53.2727 | -27.7936 | 3.343 | 53.2721 | -27.7902 | 3.344 | 288 | 0.09 | E_SP29 | 4 |
Shrawan2 | 53.0714 | -27.9353 | 3.355 | 53.0346 | -27.9438 | 3.363 | 128 | 0.93 | E_SP58 | 4 |
Shrawan3 | 53.1552 | -27.8959 | 3.242 | 53.1510 | -27.8919 | 3.237 | 342 | 0.15 | E_SP24 | 4 |
Shrawan4 | 53.2022 | -27.9406 | 3.335 | 53.1984 | -27.9349 | 3.337 | 186 | 0.18 | E_SP46 | 4 |
Smruti1 | 53.0076 | -27.7463 | 3.410 | 53.0193 | -27.7403 | 3.417 | 286 | 0.33 | E_SP30 | 4,5,6,7 |
Smruti2 | 53.0042 | -27.7411 | 3.479 | 53.0159 | -27.7498 | 3.490 | 234 | 0.37 | E_SP36 | 4,5,6,7 |
Smruti3 | 53.0613 | -27.8723 | 3.471 | 53.8723 | -27.8795 | 3.469 | 1731 | 1.96+ | E_SP1 | 4,5,6,7 |
Smruti4 | 53.2290 | -27.8828 | 3.462 | 53.8723 | -27.8795 | 3.469 | 1731 | 2.11+ | E_SP1 | 4,5,6,7 |
Smruti5 | 53.0412 | -27.7804 | 3.530 | 53.0159 | -27.7498 | 3.490 | 234 | 1.03 | E_SP36 | 4,5,6,7 |
Smruti6 | 53.1586 | -27.6964 | 3.418 | 53.1521 | -27.7078 | 3.418 | 121 | 0.25+ | E_SP60 | 4,5,6,7 |
Sparsh | 53.0579 | -27.8670 | 3.696 | 53.0478 | -27.8730 | 3.701 | 1173 | 0.29 | E_SP4 | 4,8 |
Ruchi1 | 53.2124 | -27.8306 | 4.150 | 53.2185 | -27.8273 | 4.155 | 610 | 0.16 | E_SP11 | 4 |
@@
Structure | R.A.lit | Decl.lit | R.A.C3VO | Decl.C3VO | Amp.a | dtrans | IDC3VOb | ref. | ||
@@
PClJ0227-0421.1 | 36.6375 | -4.2381 | 3.269 | 36.6969 | -4.3563 | 3.292 | 1574 | 3.69∗ | V_SP2 | 1,5 |
PClJ0227-0421.2 | 36.7438 | -4.2949 | 3.325 | 36.6969 | -4.3563 | 3.292 | 1574 | 2.14+ | V_SP2 | 1,5 |
PClJ0227-0421.3 | 36.7775 | -4.3457 | 3.299 | 36.6969 | -4.3563 | 3.292 | 1574 | 2.25+ | V_SP2 | 1,5 |
PClJ0227-0421.4 | 36.6467 | -4.3703 | 3.300 | 36.6969 | -4.3563 | 3.292 | 1574 | 1.44 | V_SP2 | 1,5 |
PClJ0227-0421.5 | 36.5817 | -4.3877 | 3.298 | 36.6969 | -4.3563 | 3.292 | 1574 | 3.31∗,+ | V_SP2 | 1,5 |
PClJ0227-0421.6 | 36.6417 | -4.4472 | 3.302 | 36.6434 | -4.4638 | 3.258 | 75 | 0.47 | V_SP76 | 1,5 |
@@ 6 Conclusions
By applying our search algorithm to our plethora of photometry and spectroscopy in the C3VO fields, we were able to find detections of 561 protostructure candidates, many of which have not previously been reported in the literature. The main takeaways from our findings are as follows:
@@
Over 400 potentially new protostructures were found in the COSMOS, ECDFS, and CFHTLS-D1 fields, all of which have some spectroscopically confirmed members.
With the aid of mock observations of simulated structures, we have found completeness and purity as high as 80% and 70%, respectively, in the best-case scenario. These numbers correspond to and structures with for completeness, and a spectral redshift fraction of % for purity (i.e., SF1.5 or better). Typical completeness and purity are around 20–40% for , though the precise value sensitively depends on the mass and redshift, with a precipitous falloff to . Purity is typically found to be 40–60%.
More than 90% of the 100 protoclusters previously reported in the literature (which were originally selected using a variety of different methods) in these three fields were recovered with our method. Of the few protoclusters that were not detected using our methodology, the vast majority were either situated near the edge of our maps or in multicomponent systems where our deblending was not aggressive enough to discriminate different subcomponents.
@@
The R.A./decl./ coordinates of all protostructure candidates are made publicly available as an electronic table on CDS. The Gaussian amplitude () of each protostructure candidate, which represents a rough mass proxy, as well as its associated error, is also provided.
Fits are provided between our mass proxy () and the mass of simulated structures for differing levels of spectroscopic completeness. These fits may be applied to the Gaussian amplitude of our candidate protostructure detections to derive crude masses for the observed candidates, though with considerable scatter and some bias. We caution that masses of our candidates can be estimated more accurately and precisely, but it must be on a case-by-case basis considering individual parameters such as the spectral sampling, tracer population, and galaxy bias for the tracers, as has been done with other C3VO studies (e.g., Cucciati et al., 2018; Shen et al., 2021; Forrest et al., 2023).
With our results, we have once again demonstrated the efficacy of the VMC mapping method and its utility when combined with any large and complete spectroscopic and photometric data set. While our work was on a fairly small region of the sky, the methodology is adaptable to any spectroscopic and photometric data set of sufficient size. This is especially important to consider in the context of active and upcoming large-scale surveys, such as those conducted by Euclid111111https://www.euclid-ec.org/, the Vera C. Rubin Observatory121212https://rubinobservatory.org/, Subaru’s Prime Focus Spectrograph131313https://pfs.ipmu.jp, the Nancy Grace Roman Space Telescope141414https://roman.gsfc.nasa.gov/, and the Maunakea Spectroscopic Explorer151515https://mse.cfht.hawaii.edu/, which are able to reach deeper magnitudes and offer larger data sets at even higher redshifts than ever before. Of these, Euclid is the only telescope currently in operation, and the Euclid Wide Survey in particular has the potential to find tens of thousands of protoclusters with masses over in its 14,500 deg2 area of coverage (Euclid Collaboration et al., 2025). Methods based largely on the VMC technique are already proving to be effective at uncovering a large number of protocluster candidates in EUCLID mock data and will be used for searches in real data (Ramos-Chernenko 2025, in preparation). Thanks to the deluge of data on the way that is focused on the redshift range that marks the transition between clusters and protoclusters, the next decade promises to be a boon for protocluster science.